# Hungarian Research Network (HUN-REN) Laboratory on Engineering and Management Intelligence Research Group of Operations Research and Decision Systems

### v1

@@ 4.1 The suboptimality of pre-assigning the host
a Skip mechanism, Qatar is pre-assigned to Group A (official draw procedure)
@@
Mexico X X X X X X X X 0.6 0.4 0.4 0.6 0.3 0.3 0.3 0.6 0.4 X 3.4 1.4 0.3 4.2 X 8.4
Netherlands X X X X X X X X 0.2 0.5 0.4 0.1 0.5 0.5 0.5 0.1 0 0 1.4 0.3 0 1.7 0 2.9
Denmark X X X X X X X X 0.1 0.4 0.5 0.2 0.5 0.5 0.4 0.1 0 0 1.4 0.3 0 1.7 0 2.9
Germany X X X X X X X X 0.1 0.4 0.5 0.1 0.5 0.5 0.5 0.1 0 0 1.4 0.3 0 1.7 0 2.9
Uruguay X X X X X X X X 1.9 3 3 1.9 1.8 1.7 3.1 1.8 0.9 0.2 X 4.1 0.8 X 0.2 2.1
Switzerland X X X X X X X X 0.2 0.5 0.4 0.2 0.5 0.5 0.5 0.1 0 0 1.4 0.3 0 1.7 0 2.9
United States X X X X X X X X 0.6 0.4 0.4 0.6 0.3 0.3 0.4 0.6 0.3 X 3.4 1.4 0.4 4.2 X 8.4
Croatia X X X X X X X X 0.1 0.5 0.5 0.1 0.5 0.5 0.4 0.1 0 0.1 1.4 0.3 0 1.7 0 2.9
Senegal 0.6 0.2 0.1 0.1 1.9 0.2 0.6 0.1 X X X X X X X X X 0.5 1.4 3 X 0 0.5 3.4
Iran 0.4 0.5 0.4 0.4 3 0.5 0.4 0.5 X X X X X X X X 0.3 0.2 3.4 X 0.3 X 0.2 2.4
Japan 0.4 0.4 0.5 0.5 3 0.4 0.4 0.5 X X X X X X X X 0.3 0.2 3.4 X 0.3 X 0.2 2.4
Morocco 0.6 0.1 0.2 0.1 1.9 0.2 0.6 0.1 X X X X X X X X X 0.5 1.4 3 X 0 0.5 3.5
Serbia 0.3 0.5 0.5 0.5 1.8 0.5 0.3 0.5 X X X X X X X X 0.5 0.4 3 4.5 0.4 0 0.5 1.5
Poland 0.3 0.5 0.5 0.5 1.7 0.5 0.3 0.5 X X X X X X X X 0.4 0.4 3 4.5 0.4 0 0.4 1.5
South Korea 0.3 0.5 0.4 0.5 3.1 0.5 0.4 0.4 X X X X X X X X 0.3 0.2 3.4 X 0.3 X 0.2 2.5
Tunisia 0.6 0.1 0.1 0.1 1.8 0.1 0.6 0.1 X X X X X X X X X 0.5 1.4 3 X 0 0.5 3.4
b Skip mechanism, the group containing Qatar is called Group A (our proposal)
@@
Mexico X X X X X X X X 0.5 0.7 0.7 0.5 0.3 0.3 0.7 0.5 0.2 X 2.8 1.8 0.1 4.2 X 8.5
Netherlands X X X X X X X X 0.2 0.3 0.3 0.2 0.3 0.3 0.4 0.2 0.1 0 1.1 0 0.1 1.7 0 3
Denmark X X X X X X X X 0.1 0.4 0.4 0.2 0.3 0.3 0.3 0.2 0.1 0 1.1 0 0.1 1.7 0 2.9
Germany X X X X X X X X 0.2 0.3 0.4 0.2 0.3 0.3 0.3 0.2 0.1 0 1.1 0 0.1 1.7 0 2.9
Uruguay X X X X X X X X 1.8 3.1 3.1 1.8 1.9 1.9 3.1 1.8 0.8 0.1 X 3.7 0.8 X 0 2.1
Switzerland X X X X X X X X 0.2 0.3 0.3 0.2 0.3 0.3 0.3 0.2 0.1 0 1.1 0 0.1 1.7 0 2.9
United States X X X X X X X X 0.5 0.7 0.7 0.5 0.3 0.2 0.7 0.5 0.2 X 2.8 1.8 0.2 4.2 X 8.4
Croatia X X X X X X X X 0.2 0.3 0.4 0.1 0.3 0.3 0.3 0.2 0.1 0 1.1 0 0.1 1.7 0 3
Senegal 0.5 0.2 0.1 0.2 1.8 0.2 0.5 0.2 X X X X X X X X X 0.4 1.4 2.2 X 0.4 0.4 2.5
Iran 0.7 0.3 0.4 0.3 3.1 0.3 0.7 0.3 X X X X X X X X 0.4 0.1 2.6 X 0.4 X 0.1 2
Japan 0.7 0.3 0.4 0.4 3.1 0.3 0.7 0.4 X X X X X X X X 0.4 0.1 2.5 X 0.3 X 0.1 2
Morocco 0.5 0.2 0.2 0.2 1.8 0.2 0.5 0.1 X X X X X X X X X 0.4 1.4 2.2 X 0.4 0.4 2.5
Serbia 0.3 0.3 0.3 0.3 1.9 0.3 0.3 0.3 X X X X X X X X 0.6 0.7 1.8 3.3 0.5 0.5 0.7 0.8
Poland 0.3 0.3 0.3 0.3 1.9 0.3 0.2 0.3 X X X X X X X X 0.6 0.7 1.8 3.3 0.6 0.5 0.7 0.8
South Korea 0.7 0.4 0.3 0.3 3.1 0.3 0.7 0.3 X X X X X X X X 0.4 0.1 2.6 X 0.4 X 0.1 2
Tunisia 0.5 0.2 0.2 0.2 1.8 0.2 0.5 0.2 X X X X X X X X X 0.3 1.3 2.2 X 0.3 0.3 2.5
@@
These results are reinforced by Table 4, which compares the distortions for all team pairs under the two versions of the Skip mechanism. The advantage of our proposal is the most visible for the host Qatar, where the greatest biases—playing against Senegal, Morocco, Serbia, Poland, Tunisia, and the winner of UEFA PO—substantially decrease. Labelling the groups after the draw also makes the Skip mechanism less unfair for Serbia, Poland (the two European teams in Pot 3), and Ecuador.
Consequently, we have a clear suggestion for improving the fairness of the 2022 FIFA World Cup draw by not assigning the host Qatar to Group A automatically.
4.2 The impact of draw constraints
00222244446666888810101010121212121414141416161616181818182020202022222222242424242626262628282828303030300.10.10.10.1111110101010100100100100Set of constraints imposed (see Table 3)Proportion of valid assignments (%, log scale)Ratio of feasible solutions without Constraint E                       Ratio of feasible solutions with Constraint E
Figure 2: The effects of draw constraints and mechanisms:
the chance of feasibility and relative inequality distortion
Figure 2 shows the proportion of valid assignments under the 32 sets of draw constraints. The most restrictive is Constraint E, affecting the UEFA teams, which reduces the probability of feasibility by at least 79%. On the other hand, Constraint C for the CONCACAF nations decreases this chance only between 34% and 46%. Imposing all restrictions means that only one out of 532 unconstrained draws can be accepted on average.
a Measure of non-uniformity: Mean absolute bias $\Delta(\mathcal{C})$
0.10.10.10.1111110101010100100100100000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.4Ratio of feasible solutions (%, log scale)Mean absolute bias (percentage point)
0.10.10.10.11111101010101001001001000022224444666688881010101012121212Ratio of feasible solutions (%, log scale)Maximum absolute bias (percentage point)Skip mechanism without Constraint E                                              Skip mechanism with Constraint E                                                  Skip mechanism with all constraints if Qatar is pre-assigned to Group A
Figure 3: The aggregated bias of the Skip mechanism for the 32 sets of draw constraints
Figure 3 presents the distortion of the draw under all sets of draw constraints as a function of the proportion of valid assignments. Imposing more constraints tends to increase the bias of the Skip mechanism but there is no functional relationship between the restrictiveness of the constraints and the implied bias. The draw procedure clearly struggles with Constraint E. In the absence of this restriction, $\Delta(\mathcal{C})$ ($\Omega(\mathcal{C})$) never exceeds 0.72 (3.1), but $\Delta(\mathcal{C})$ ($\Omega(\mathcal{C})$) is certainly higher than 0.97 (6.2) if Constraint E is imposed. Constraint E is more challenging for the Skip mechanism since it allows more than one team from a given set to be assigned to the same group (Csató, 2025c, ). On the other hand, Constraints A–D can be guaranteed by prohibited pairs, thus, imposing only one of them preserves uniformity for the Skip mechanism (Csató, 2025c, , Lemma 1). These results are insensitive to the measure of non-uniformity, the patterns seen in Figures 3.a and 3.b are similar.
4.3 Identifying draw constraints on the Pareto frontier
a Measure of non-uniformity: Mean absolute bias $\Delta(\mathcal{C})$
55555.55.55.55.566666.56.56.56.577777.57.57.57.588888.58.58.58.599999.59.59.59.51010101010.510.510.510.5000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.4Average number of intra-continental group matchesMean absolute bias (percentage point)
55555.55.55.55.566666.56.56.56.577777.57.57.57.588888.58.58.58.599999.59.59.59.51010101010.510.510.510.50022224444666688881010101012121212Average number of intra-continental group matchesMaximum absolute bias (percentage point)Skip mechanism without Constraint E      Skip mechanism with Constraint E
Figure 4: The trade-off between intra-continental group matches and non-uniformity
Note: The dotted black line shows the Pareto frontier.
Figure 4 uncovers the trade-off between intra-continental group matches and the non-uniformity of the draw. The expected number of unattractive games is 10.1 in an unconstrained draw that can be decreased to 8.8 without any price by imposing Constraint A. The Pareto frontier also contains the draw under Constraints A–D, which implies about 6.5 intra-continental group matches for a moderate level of non-uniformity. Finally, imposing all restrictions is also efficient by fixing the number of intra-continental games at five, albeit its price may be excessive.
Table 5: The optimal sets of draw constraints
Problem Measure of non-uniformity Optimal $\mathcal{C}$ Value of $\alpha$
(3) $\Delta(\mathcal{C})$ 16 (A) $0.2837\leq\alpha$
(3) $\Delta(\mathcal{C})$ 18 (A, D) $0.2293\leq\alpha\leq 0.2837$
(3) $\Delta(\mathcal{C})$ 30 (A–D) $0.2105\leq\alpha\leq 0.2293$
(3) $\Delta(\mathcal{C})$ 31 (A–E) $\alpha\leq 0.2105$
(4) $\Omega(\mathcal{C})$ 16 (A) $0.436\leq\alpha$
(4) $\Omega(\mathcal{C})$ 30 (A–D) $0.2127\leq\alpha\leq 0.436$
(4) $\Omega(\mathcal{C})$ 31 (A–E) $\alpha\leq 0.2127$
Table 5 summarises the potential optimal solutions to problems (3) and (4), respectively. Besides the three sets of constraints discussed above, only one set, Constraints A and D, is efficient if the non-uniformity measure is mean absolute bias $\Delta(\mathcal{C})$. For problem (3), scenario 30 with Constraints A–D is optimal only in a narrow range of parameter $\alpha$.
Therefore, while the official policy of imposing all geographic constraints can obviously be rationalised if one wants to minimise the number intra-continental group games, the results of our analysis remain interesting for the organiser: some further sets of draw restrictions might also be chosen if the non-uniformity of the draw is judged excessive and the stakeholders, especially the teams, call for less unfair rules.
5 Concluding remarks
Our proposal is applied to the 2022 FIFA World Cup draw with the following findings:
  • The unnecessary policy of pre-assigning Qatar increases the mean (maximum) absolute bias $\Delta(\mathcal{C})$ ($\Omega(\mathcal{C})$) of the draw by 0.17 (2) percentage points (Section 4.1);
  • The Skip mechanism struggles with Constraint E concerning UEFA teams, which increases the mean (maximum) absolute bias $\Delta(\mathcal{C})$ ($\Omega(\mathcal{C})$) of the draw by at least 0.44 (4.9) percentage points (Section 4.2);
  • Out of 32 reasonable scenarios, at most four sets of draw constraints provide an optimal trade-off between the expected number of intra-continental group matches and the non-uniformity of the Skip mechanism (Section 4.3).
The suboptimality of draw procedure for the 2022 FIFA World Cup has been revealed. Consequently, tournament organisers are encouraged to revise the pre-assignment of the host(s), as well as to carry out similar investigations before the group draw of major competitions in order to compute the efficient choices with respect to the inevitable trade-off between the attractiveness of group matches and the fairness of the draw.
Acknowledgements
This paper could not have been written without my father (also called László Csató), who has helped to code the simulations in Python.

### v2

@@ Acknowledgements
a Skip mechanism, the host (Russia) is pre-assigned to Group A (official draw procedure)
Spain
Peru
Switzerland
England
Colombia
Mexico
Uruguay
Croatia
Denmark
Iceland
Costa Rica
Sweden
Tunisia
Egypt
Senegal
Iran
Serbia
Nigeria
Australia
Japan
Morocco
Panama
South Korea
Saudi Arabia
Russia 3.3 4.2 3.3 3.3 4.2 0.8 4.2 3.3 4.3 4.3 1.2 4.3 2.9 2.9 2.9 3 4.7 2.1 1.8 1.7 2.1 1.8 1.8 1.8
Germany 0.3 0.8 0.3 0.3 0.8 3.8 0.9 0.3 0.5 0.5 0.1 0.5 0.4 0.4 0.4 0.3 1.1 0.4 0 0 0.4 0.3 0 0
Brazil 2.4 2.4 2.5 9.8 2.5 0.9 1 1 1 0.4 0.4 0.4 0.7 5.2 0 0.8 0.9 0 1.7 0.9 0.9
Portugal 0.3 0.8 0.3 0.3 0.8 3.8 0.8 0.3 0.5 0.5 0.2 0.5 0.4 0.4 0.4 0.3 1.1 0.4 0 0 0.4 0.3 0 0
Argentina 2.4 2.5 2.5 9.9 2.5 1 0.9 1 1 0.4 0.4 0.4 0.7 5.2 0 0.9 0.9 0 1.6 0.9 0.9
Belgium 0.3 0.8 0.3 0.3 0.8 3.7 0.8 0.3 0.5 0.5 0.2 0.5 0.4 0.4 0.4 0.3 1.2 0.4 0 0 0.5 0.3 0 0
Poland 0.3 0.8 0.3 0.3 0.9 3.8 0.8 0.4 0.5 0.5 0.2 0.5 0.4 0.4 0.4 0.3 1.1 0.4 0 0 0.4 0.3 0 0
France 0.3 0.8 0.4 0.4 0.8 3.8 0.8 0.3 0.5 0.5 0.2 0.5 0.4 0.4 0.4 0.3 1.2 0.4 0 0 0.4 0.3 0 0
Spain 0.3 0.4 0.3 0.3 0.3 0.3 0.3 0.5 1.1 0.1 0.2 0.2 0.1 0.2 0.2 0.2
Peru 0.6 0.6 0.4 0.6 0.5 0.5 0.5 0.8 2 0.5 0.3 0.3 0.5 0.2 0.3 0.3
Switzerland 0.3 0.4 0.2 0.3 0.2 0.3 0.3 0.5 1.1 0.1 0.2 0.2 0.1 0.2 0.2 0.2
England 0.3 0.4 0.3 0.3 0.3 0.2 0.3 0.5 1.1 0.1 0.2 0.2 0.1 0.2 0.2 0.2
Colombia 0.6 0.6 0.4 0.6 0.5 0.5 0.5 0.8 1.9 0.5 0.3 0.3 0.5 0.2 0.3 0.3
Mexico 0.5 0.4 0.5 0.4 0.4 0.4 0.2 10.2 1.7 1.7 1.7 1.7 1.7 1.7
Uruguay 0.6 0.6 0.4 0.6 0.5 0.5 0.5 0.8 1.9 0.5 0.3 0.3 0.4 0.2 0.3 0.3
Croatia 0.4 0.3 0.3 0.4 0.3 0.3 0.3 0.5 1.1 0.1 0.2 0.2 0.1 0.2 0.2 0.2
Denmark 0.3 0.6 0.3 0.3 0.6 0.5 0.6 0.4 1.7 0.1 0.3 0.3 0.1 0.7 0.3 0.3
Iceland 0.4 0.6 0.4 0.4 0.6 0.4 0.6 0.3 1.7 0.1 0.3 0.3 0.1 0.7 0.3 0.3
Costa Rica 0.3 0.4 0.2 0.3 0.4 0.4 0.3 0.5 0.3 0.2 0.3 0.2 0.3 0.3
Sweden 0.3 0.6 0.3 0.3 0.6 0.5 0.6 0.4 1.7 0.1 0.3 0.3 0.1 0.7 0.3 0.3
Tunisia 0.3 0.5 0.2 0.3 0.5 0.4 0.5 0.3 0.8 0.2 0.2 0.1 0.2 0.2
Egypt 0.3 0.5 0.3 0.2 0.5 0.4 0.5 0.3 0.8 0.2 0.2 0.1 0.2 0.2
Senegal 0.3 0.5 0.3 0.3 0.5 0.4 0.5 0.3 0.8 0.2 0.3 0.1 0.2 0.2
Iran 0.5 0.8 0.5 0.5 0.8 0.2 0.8 0.5 2.2 0.1 0.1 2.4
b Skip mechanism, the group with the host (Russia) is called Group A (our proposal)
Spain
Peru
Switzerland
England
Colombia
Mexico
Uruguay
Croatia
Denmark
Iceland
Costa Rica
Sweden
Tunisia
Egypt
Senegal
Iran
Serbia
Nigeria
Australia
Japan
Morocco
Panama
South Korea
Saudi Arabia
Russia 0.8 0 0.8 0.8 0 3.3 0 0.8 0.1 0.1 0.3 0.1 0 0 0 0.1 1.7 0.1 0.2 0.2 0.2 0.5 0.2 0.3
Germany 0.9 0 0.8 0.8 0 3.3 0 0.8 0.1 0.1 0.3 0.1 0 0 0 0.1 1.6 0.1 0.3 0.2 0.1 0.5 0.2 0.2
Brazil 2.5 2.5 2.5 9.9 2.5 0.4 0.4 0.8 0.4 0 0 0 0.3 4.9 0.4 0.7 0.7 0.4 1.5 0.7 0.7
Portugal 0.8 0 0.8 0.8 0 3.3 0 0.8 0.1 0.1 0.3 0.1 0 0 0 0.1 1.6 0.1 0.2 0.2 0.1 0.5 0.2 0.2
Argentina 2.5 2.5 2.4 9.8 2.4 0.3 0.4 0.8 0.4 0 0 0 0.3 5 0.4 0.7 0.7 0.4 1.5 0.7 0.7
Belgium 0.8 0 0.8 0.8 0 3.3 0 0.8 0.1 0.1 0.3 0.1 0 0 0 0.1 1.7 0.1 0.2 0.2 0.1 0.5 0.2 0.2
Poland 0.8 0 0.8 0.8 0 3.3 0 0.8 0.1 0.1 0.3 0.1 0 0 0 0.1 1.7 0.2 0.2 0.2 0.1 0.5 0.2 0.2
France 0.8 0 0.8 0.8 0 3.3 0 0.8 0.1 0.2 0.3 0.1 0 0 0 0.1 1.6 0.1 0.2 0.2 0.1 0.5 0.2 0.2
Spain 0.6 0.6 0.1 0.6 0.4 0.4 0.4 0.7 1.3 0 0.3 0.3 0 0.2 0.3 0.3
Peru 0.9 0.9 0.2 0.9 0.7 0.7 0.7 1 1.8 0.6 0.2 0.3 0.6 0.3 0.2 0.2
Switzerland 0.6 0.6 0.2 0.6 0.4 0.4 0.4 0.7 1.3 0 0.2 0.3 0 0.3 0.3 0.2
England 0.6 0.6 0.2 0.6 0.4 0.4 0.4 0.7 1.3 0 0.3 0.3 0 0.3 0.3 0.3
Colombia 1 0.9 0.2 0.9 0.7 0.7 0.6 1 1.8 0.6 0.2 0.2 0.6 0.3 0.3 0.2
Mexico 0.5 0.4 0.5 0.4 0.4 0.4 0.3 10.7 1.7 1.8 1.8 1.7 1.8 1.8
Uruguay 0.9 0.9 0.2 0.9 0.7 0.7 0.7 1 1.8 0.6 0.2 0.2 0.6 0.3 0.2 0.2
Croatia 0.6 0.6 0.2 0.6 0.4 0.4 0.4 0.7 1.3 0 0.3 0.3 0 0.2 0.3 0.3
Denmark 0.6 0.9 0.6 0.6 1 0.5 0.9 0.6 1.5 0.1 0.3 0.3 0.1 0.7 0.2 0.3
Iceland 0.6 0.9 0.6 0.6 0.9 0.4 0.9 0.6 1.5 0.2 0.3 0.2 0.1 0.7 0.3 0.3
Costa Rica 0.1 0.2 0.2 0.2 0.2 0.2 0.2 0.7 0.1 0.3 0.3 0.2 0.2 0.2
Sweden 0.6 0.9 0.6 0.6 0.9 0.5 0.9 0.6 1.5 0.1 0.3 0.2 0.1 0.6 0.3 0.3
Tunisia 0.4 0.7 0.4 0.4 0.7 0.4 0.7 0.4 0.6 0.2 0.2 0.1 0.2 0.2
Egypt 0.4 0.7 0.4 0.4 0.7 0.4 0.7 0.4 0.6 0.2 0.2 0.1 0.2 0.2
Senegal 0.4 0.7 0.4 0.4 0.6 0.4 0.7 0.4 0.6 0.2 0.2 0.1 0.2 0.2
Iran 0.7 1 0.7 0.7 1 0.3 1 0.7 1.9 0.2 0.2 2.3
  • The probabilities are based on 10 million simulated draws.
  • Blank cells represent pairs of teams that cannot play in the group stage.
  • Green (Red) colour means that the Skip mechanism implies a higher (lower) probability than a uniform draw. For instance, the probability of assigning Russia and Nigeria to the same group is lower (higher) by about 2.1 (0.1) percentage points with the official draw mechanism (our proposal) compared to a uniform draw, when it is 12.5%.
  • Darker colour indicates a higher value.
Table A.2: The non-uniformity of the 2022 FIFA World Cup draw
a Skip mechanism, the host (Qatar) is pre-assigned to Group A (official draw procedure)
@@
Mexico 0.6 0.4 0.4 0.6 0.3 0.3 0.3 0.6 0.4 3.4 1.4 0.3 4.2 8.4
Netherlands 0.2 0.5 0.4 0.1 0.5 0.5 0.5 0.1 0 0 1.4 0.3 0 1.7 0 2.9
Denmark 0.1 0.4 0.5 0.2 0.5 0.5 0.4 0.1 0 0 1.4 0.3 0 1.7 0 2.9
Germany 0.1 0.4 0.5 0.1 0.5 0.5 0.5 0.1 0 0 1.4 0.3 0 1.7 0 2.9
Uruguay 1.9 3 3 1.9 1.8 1.7 3.1 1.8 0.9 0.2 4.1 0.8 0.2 2.1
Switzerland 0.2 0.5 0.4 0.2 0.5 0.5 0.5 0.1 0 0 1.4 0.3 0 1.7 0 2.9
United States 0.6 0.4 0.4 0.6 0.3 0.3 0.4 0.6 0.3 3.4 1.4 0.4 4.2 8.4
Croatia 0.1 0.5 0.5 0.1 0.5 0.5 0.4 0.1 0 0.1 1.4 0.3 0 1.7 0 2.9
Senegal 0.6 0.2 0.1 0.1 1.9 0.2 0.6 0.1 0.5 1.4 3 0 0.5 3.4
Iran 0.4 0.5 0.4 0.4 3 0.5 0.4 0.5 0.3 0.2 3.4 0.3 0.2 2.4
Japan 0.4 0.4 0.5 0.5 3 0.4 0.4 0.5 0.3 0.2 3.4 0.3 0.2 2.4
Morocco 0.6 0.1 0.2 0.1 1.9 0.2 0.6 0.1 0.5 1.4 3 0 0.5 3.5
Serbia 0.3 0.5 0.5 0.5 1.8 0.5 0.3 0.5 0.5 0.4 3 4.5 0.4 0 0.5 1.5
Poland 0.3 0.5 0.5 0.5 1.7 0.5 0.3 0.5 0.4 0.4 3 4.5 0.4 0 0.4 1.5
South Korea 0.3 0.5 0.4 0.5 3.1 0.5 0.4 0.4 0.3 0.2 3.4 0.3 0.2 2.5
Tunisia 0.6 0.1 0.1 0.1 1.8 0.1 0.6 0.1 0.5 1.4 3 0 0.5 3.4
b Skip mechanism, the group with the host (Qatar) is called Group A (our proposal)
@@
Mexico 0.5 0.7 0.7 0.5 0.3 0.3 0.7 0.5 0.2 2.8 1.8 0.1 4.2 8.5
Netherlands 0.2 0.3 0.3 0.2 0.3 0.3 0.4 0.2 0.1 0 1.1 0 0.1 1.7 0 3
Denmark 0.1 0.4 0.4 0.2 0.3 0.3 0.3 0.2 0.1 0 1.1 0 0.1 1.7 0 2.9
Germany 0.2 0.3 0.4 0.2 0.3 0.3 0.3 0.2 0.1 0 1.1 0 0.1 1.7 0 2.9
Uruguay 1.8 3.1 3.1 1.8 1.9 1.9 3.1 1.8 0.8 0.1 3.7 0.8 0 2.1
Switzerland 0.2 0.3 0.3 0.2 0.3 0.3 0.3 0.2 0.1 0 1.1 0 0.1 1.7 0 2.9
United States 0.5 0.7 0.7 0.5 0.3 0.2 0.7 0.5 0.2 2.8 1.8 0.2 4.2 8.4
Croatia 0.2 0.3 0.4 0.1 0.3 0.3 0.3 0.2 0.1 0 1.1 0 0.1 1.7 0 3
Senegal 0.5 0.2 0.1 0.2 1.8 0.2 0.5 0.2 0.4 1.4 2.2 0.4 0.4 2.5
Iran 0.7 0.3 0.4 0.3 3.1 0.3 0.7 0.3 0.4 0.1 2.6 0.4 0.1 2
Japan 0.7 0.3 0.4 0.4 3.1 0.3 0.7 0.4 0.4 0.1 2.5 0.3 0.1 2
Morocco 0.5 0.2 0.2 0.2 1.8 0.2 0.5 0.1 0.4 1.4 2.2 0.4 0.4 2.5
Serbia 0.3 0.3 0.3 0.3 1.9 0.3 0.3 0.3 0.6 0.7 1.8 3.3 0.5 0.5 0.7 0.8
Poland 0.3 0.3 0.3 0.3 1.9 0.3 0.2 0.3 0.6 0.7 1.8 3.3 0.6 0.5 0.7 0.8
South Korea 0.7 0.4 0.3 0.3 3.1 0.3 0.7 0.3 0.4 0.1 2.6 0.4 0.1 2
Tunisia 0.5 0.2 0.2 0.2 1.8 0.2 0.5 0.2 0.3 1.3 2.2 0.3 0.3 2.5